Systems Staff Meeting Notes 2/22/99

Systems Staff Meeting Notes
February 22, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Announcements
  1. Security Scan - two of our networks were scanned last Thursday, as scheduled. Subnet 10 was missed - it was supposed to be done on Friday, but the network was down. They finally scanned it this weekend. The first scan did generate a 29 paged report that listed six high risk factors. It indicated risk with the DNS server and the SNMP public print box. The auditor could not read the final report, but assured us that NERDC could give reference. NERDC said it would be difficult to provide. We did change the community string on 3 of 4 printers - the fourth one can only be accessed via Telnet. Our vulnerability is about as low as it can get. The subnet 10 scan showed the Back Orafice station opened (we left it that way) and 4 or 5 printers at risk. There is some question as to how effective these scans are, there was a server on campus that was scanned, fixed then hacked into. Another department had their mail server crash because of the insistence to run the scan at prime time. In the future, we may want to resist these scans.
  2. AIX box - we shut down SEND MAIL and found that this caused forms to stop working. We thought using cookies would bypass it, but that was not the case. We needed some mail way to send forms - NERDC suggested shutting off the incoming mail & keeping the outgoing. This was successful once Caroline was rebooted.
  3. ATM switches - NERDC has decided to reboot them every few weeks - this straightens out potential problems that cause the switches to get confused - as happened last week. Larger packets cause more problems then small ones. This outage produced a call to keep the SNA terminals, given that they were providing the only access to NOTIS & LUIS.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Installed a switch in the 1st floor Com closet to help with the phantom problems in Preservation. This CISCO switch is not at the current software level, we need to update it.
Liaisons
  1. LMG this week will address the enhanced Liaison Training Program. The higher certification phase was put on hold by the Directors. The workshop with out-of-town speakers is still a go.
Networking
  1. Estimates for the upcoming wiring projects have still not been received.
Software Projects Status
  1. IR database - Ying is working on the basic function - figuring out what tables to retain, modifications to make - she will make a list for Bill to review. Test data from Fox to SQL server is about 1/3 complete - there's a lot of data cleanup to do, the indexes are pretty corrupted. SQL server will handle these discrepancies better.
  2. Resource Services - licensing - Martha has an interest in getting it out to them. Brian has noticed a problem since the switch from NT2 to NT7 - access device path is denied - it will not allow Resource Services to access the database on NT7, even though the permissions look right. Brian has narrowed it down to the menus - license, requests & copies will be ready soon. Rich should be able to use it to work with his IP's.
  3. WebPaula & the Trouble Ticket Viewer will both be accessed via the web.
  4. Special Collections database is finished.
  5. Mellon demo of latest incarnation scheduled for this afternoon - Bill & Winston will meet with Martha Hruska, Erich Kesse & Cecilia Botero.
DP Coordinator
  1. Peter McKay has talked to Compustat, they will try shipping another test shipment of tapes in ASCII (rather than EBCDIC) format. This will require access & processing change. Peter inquired about CD ROM format, but proprietary type software is required.
Calendar
  1. St. Patrick's Day - Wednesday, March 17.


Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu